• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name pawel

    Log entry time 08:01:15 on November16,2002

    Entry number 90173

    keyword=Owl Shift Summary (Nov. 15 & 16)

    Owl Shift Crew: Mark Jones, Pawel Ambrozewicz

    Nov. 15 0:00 - 8:00:

    1:50 MCC calls. Hall C requested Pockel Cell OFF.
    6:35 Communication lost with ROC's and TS's (run # 3184). Rebooting crates from
    SLOW CONTROLS GUI and resetting CODA restores the communication.

    Production running:
    Run Events Sum-trues (for all runs up to listed run)
    3176 1.0M 28
    3177 1.0M 30
    3178 1.0M 32
    3179 0.7M 34
    3180 1.0M 34
    3181 1.0M 36
    3182 1.0M 36
    3183 1.0M 36
    3185 1.0M 37


    Nov. 16 0:00 - 8:00:
    1:50 DAQ crashed. After restart, all back to normal.
    2:23 MCC called. No beam for couple of minutes due to FSD problem.
    2:38 MCC said it would deliver the beam back.
    5:00 FPP HV turned off.
    6:20 VDC gas flow shut down.
    6:45 Gas flow restored.